On a river, a kayaker travels 2 miles upstream and 2 miles downstream in a total of 5 hours. In still water, the kayaker can travel at an average speed of 2 mph. Based on this information, what is the average speed of the current of this river?
Answer by mananth(16946)

let speed of current be x mph
against current speed = 2-x
with current speed = 2+x
distance = 2miles
time upstream + time down stream = 5
t=d/r
2/(2-x)+ 2/(2+x)=5
LCD = (2+x)(2-x)
multiply by the equation by the LCD
2(2+x)+2(2-x)= 5(4-x^2)
4+2x+4-2x= 20-5x^2
8=20-5x^2
5x^2= 20-8
5x^2=12
x^2= 12/5
x= sqrt(12/5)
x= 1.55 mph speed of current